OpenStudy (anonymous):

Challenge! Ready? Given the proportion, a/b=8/15, what ratio completes the equivalent proportion a/8=? Show me your work so that I, Lolly, know that you are not cheating.

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

you cross multiply

OpenStudy (danjs):

\[\frac{ a }{ b }=\frac{ 8 }{ 15 }\]

OpenStudy (danjs):

15a = 8b \[\frac{ 15a }{ 8 } = b\]

OpenStudy (danjs):

can you finish,, divide both sides by 15

OpenStudy (anonymous):

IT EQUALS a / 8 = b / 15

OpenStudy (anonymous):

right guys?

OpenStudy (danjs):

yes